Q1: Use the parallelogram of forces method to find the
magnitude and direction of the resultant of a force of 20N
acting at an angle of 156° and a force of 34N acting at an
angle of 243o Q2: Determine the magnitude and direction of the resultant
of a force of 15N acting horizontally to the right and a force
of 20N, inclined at an angle of 60° to the 15N force. Use
the triangle of forces method.
1
Expert's answer
2019-12-13T10:11:32-0500



Using theorem of "cos"

"F^2=20^2+34^2-2\cdot 20\cdot 34\cdot cos(93)"

"F=40 N"

The angle of direction

"\alpha=213"



Using theorem of "cos"

"F^2=20^2+15^2-2\cdot 20\cdot 15\cdot cos(120)"

"F=30 N"

The angle of direction

"\alpha=34"